Solve the system by substitution.
3x + 4y = 6
y = 2x – 4

Answer by checkley75(3666)

Y=2X-4 NOW SUBSTITUTE (2X-4) FOR Y IN THE OTHER EQUATION.
3X+4(2X-4)=6
3X+8X-16=6
11X=6+16
11X=22
X=22/11
X=2 ANSWER.
Y=2*2-4
Y=4-4
Y=0 ANSWER.
PROOF
3*2+4*0=6
6=6